What is the displacement of spring with a spring constant of 250 N/m if the restoring force is 85 N?
1
Expert's answer
2012-10-19T08:56:20-0400
According to the Hooke's law:

F = k*ΔL ==> ΔL = F/k = 85[N]/250[N/m] = 0.34 m.

So, the displacement of spring is 34 cm.